From mar@roe.ac.uk Wed Feb 29 18:00:45 2012 Date: Wed, 29 Feb 2012 14:37:02 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: VDFS meeting minutes 24/2/12 Present: ETWS, RPB, NJC, ACD, MAR, MSH, RSC, RGM, NCH Apologies: JPE, AL, KTN, STV, C37, 89X DoNM: 10am Friday March 9th 2012 in the VH (NCH in the chair) Actions discharged this time: ----------------------------- ACTION: MAR to check that non-survey flat-file access for VISTA PI programmes is all in hand. Seems to be working, at least it does for n088a0416a. Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB to roll out latest curation client / DB server share configuration for routine ops use. Continues; RSC reported that a software branch that uses the new configuration has been created. The shares still need to be set-up. ACTION: ETWS to sort out common SQL function definition script(s) and ensure schema browser parser works correctly with those. RSC noted that a common schema file had been created. An update to the schema browser parser is still required. Actions carried forward from 10/02/12 meeting --------------------------------------------- ACTION: RPB to set up WSA mirror/sync for flat file DBs using MS SQL replication. Continues; a list of tables to publish was agreed. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues. ACTION: RPB to complete the LTO-4 tape backups of VVV and SSA DBs and SSS flat file datasets. Continues; a linux client has been installed and the SSS backed up. Specific points and new actions: -------------------------------- Project management: Nothing for the minutes. WFCAM, VISTA and VST updates: No news to report this time. Comments and issues arising from CASU minutes: No new minutes this time. NCH noted that JRF's algorithm for flagging areas affected by bright stars is not to be incorporated into the CASU pipeline. Networking: All fine on the networking front. WSA/VSA/OSA Operations: ETWS reported: -- Updated 3rd party software on thoth, which concludes the update on all major curation servers. -- WFCAM: Transferred WFCAM data up to December 2011 and ingested all the metadata and catalogue data up to December 2011 as well. -- GES: Worked on the schemas with ACD and included GES parsing into the BrowserParser script. ETWS asked that as we appeared to be transferring VST data from surveys other than Atlas should we be doing anything with them. It was agreed that we might as well set up programmes in case they turn out to be needed. Hardware and Systems: Kakai and Ramses12 are back from Eclispe and appear to be working. MSH reported that orders are in for a new cluster node and a new processing machine. Currently we cannot afford to upgrade the disks on Ramses9. Software: RSC reported: Primarily been putting the finishing touches to development work on migrating to the dedicated ingest/outgest disks to manage disk space better (takes storage space pressure off of the NAS boxes and catalogue servers and guarantees evenly loaded volumes) and improve performance (these disks could in future even be replaced with high-speed RAID arrays configured without redundancy for further performance increases as they only contain temporary scratch files). Created dev_diskperf2 branch to allow operations to continue testing parallel CU3 & 4 tasks before final live roll out. Also, a VMCDR1 release database has been prepared using the latest version of CU19 to exclude fields not contained within the ESO release. This will be ready for release at the agreed time, subject to successful further tests of the contents being correct. NJC reported he had made a few modifications to speed up CU6. ETWS had been working on the CU3 C++ code for Atlas. Survey Data Release: VISTA public release update: VVV: still processing; ETA late March / early April VMC: Maria-Rosa has agreed to a public release to coincide with NAM. VIDEO: Still some uncertainty over exact contents. VIKING: Still waiting to hear back from Will over field selection/ deprecations. VHS: Following Richard McMahon's approval DR1 has been released. ACTION: NJC to contact Matt Jarvis and Will Sutherland to try and finalise the VIDEO and VIKING public release contents. It was suggested that the VSA webpages should include more details of the public releases: coverage, stats etc. ACTION: MAR to update VSA pages with details of public survey releases. WSA : GPS has been added to UKIDSS DR8. Non-survey Data Release: U/09B/8 was currently being prepared for release. Astrogrid deployment & Data Analysis services: MSH reported that VHS DR1 TAP and SIAP services had been set-up and that he was working on the coverage maps for Aladin. MSH has not yet been given the OK to migrate all the Astrogrid services to WFAU's servers. Miscellaneous: RGM mentioned that the Astronomy Computing Journal was going ahead. -- Scanned by iCritical.